OpenStudy (unklerhaukus):

\[y^{\prime\prime}+4y^\prime+4y=16xe^{-2x}\]

OpenStudy (unklerhaukus):

\[y_c^{\prime\prime}+4y_c^\prime+4y_c=0\] \[m^2+4m+4=0\]\[(m+2)^2=0\]\[m=-2\] \[y_c=(A+Bx)e^{-2x}\]

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+4y_p^\prime+4y_p=16xe^{-2x}\] \[y_p=Cx^3e^{-2x}\]\[y_p^\prime= C(3x^2-2x^3)e^{-2x}\]\[y_p^{\prime\prime}=C(6x-6x^2-2(3x^2-2x^3))e^{-2x}=C(6x-12x^2+4x^3)e^{-2x}\] \[C(6x-12x^2+4x^3)e^{-2x}+4(C(3x^2-2x^3)e^{-2x})+4(Cx^3e^{-2x})=16e^{-2x}\]\[C(6x-12x^2+4x^3+12x^2-8x^3+4x^3)e^{-2x}=16e^{-2x}\]\[6Cxe^{-2x}=16e^{-2x}\]\[C=\frac 83\] \[y_p=\frac38x^3e^{-2x}\]

OpenStudy (goformit100):

nice @UnkleRhaukus

OpenStudy (anonymous):

no no why is your 8/3 suddenly turned to 3/8 lol

OpenStudy (unklerhaukus):

ok so have got the right answer, but i dont understand why the trial for the particular solution was assumed to be of this form\[y_p=Cx^3e^{-2x}\]

OpenStudy (unklerhaukus):

oh yeah i did not mean to invert the fraction for no reason

OpenStudy (unklerhaukus):

\[y_p=\frac83x^3e^{-2x}\]\[y=y_c+y_p\]\[y(x)=(A+Bx)e^{-2x}+\frac83x^3e^{-2x}\]\[y(x)=(A+Bx+\frac83x^3)e^{-2x}\]

OpenStudy (unklerhaukus):

how come \[y_p=Cx^2e^{-2x}\] didn't work?

OpenStudy (anonymous):

\[16xe^{-2x}\] based on the table of undetermined coef \[y_p = (Ax+B)e^{-2x}\] if we time only x \[Bxe^{-2x}\] will be the same with one of the homogenous solution so we time x^2 \[y_p = x^2(Ax+B)e^{-2x}\]

OpenStudy (unklerhaukus):

wh not this?\[y_p=x(C+Dx)e^{-2x}\]

OpenStudy (anonymous):

Cx e^(-2x) will be the same with the homogenous solution Bx e^(-2x)

OpenStudy (unklerhaukus):

dosent the Dx^2e^{-2x} have an extra factor of x though/?

OpenStudy (anonymous):

^ doesn't matter tho o-o.., as long as it's not the same with the homogenous solution Ae^(-2x) + Bx e^(-2x)

OpenStudy (unklerhaukus):

i dont get it

OpenStudy (goformit100):

yes

OpenStudy (unklerhaukus):

is there some way we can know \[y_p = x^2(C+Dx)e^{-2x}\] the C will be zero?

OpenStudy (anonymous):

i don't know about that ._.

OpenStudy (unklerhaukus):

when there is two constants the differentiation becomes very confusing

OpenStudy (unklerhaukus):

so i should really go back and use\[y_p = x^2(C+Dx)e^{-2x}\] and find \(y^\prime_p,\quad y_p^{\prime\prime}\) again and do the substitution again...

OpenStudy (anonymous):

i check, your answer is correct tho o-o.. C probably will get cancelled out. but i don't know how to tell if C or D will get cancelled, tho.

OpenStudy (unklerhaukus):

\[y_p=(C+Dx)x^2e^{-2x}\] \[y_p^\prime= (Dx^2+2(C+Dx)x-2(C+Dx)x^2)e^{-2x}\]\[\qquad= (2Cx-2Cx^2+3Dx^3-2Dx^3)e^{-2x}\]\[\qquad\qquad= \left(2C(x-x^2)+D(3x^3-2x^3)\right)e^{-2x}\]

OpenStudy (anonymous):

i got the same so far :P

OpenStudy (goformit100):

me too

OpenStudy (unklerhaukus):

*\[\qquad\qquad= \left(2C(x-x^2)+D(3x^2-2x^3)\right)e^{-2x}\]

OpenStudy (anonymous):

ok.. the small 2 and 3 are very hard to different -_- sorry

OpenStudy (anonymous):

here for you to check\[y_p''=(2C-8Cx+4Cx^2+6Dx-12Dx^2+4Dx^3)e^{-2x}\]

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}=\left(2C(1-2x)+D(6x-6x^2)-2\left(2C(x-x^2)+D(3x^2-2x^3)\right)\right)e^{-2x}\]\[=\left(2C-4Cx+6Dx-6Dx^2-4Cx+4Cx^2+6Dx^2-4Dx^3\right)e^{-2x}\]\[=\left(2C-8Cx+4Cx^2+6Dx-12Dx^2-4Dx^3\right)e^{-2x}\]

OpenStudy (unklerhaukus):

oh goodie i have what you have.. now i have substitute and solve for C,D

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+4y_p^\prime+4y_p=16xe^{-2x}\] \[y_p=(C+Dx)x^2e^{-2x}\]\[y_p^{\prime}=\left(2C(x-x^2)+D(3x^2-2x^3)\right)e^{-2x}\]\[y_p^{\prime\prime}=\left(C(2-8x+4x^2)+D(6x-12x^2-4x^3)\right)e^{-2x}\] \[\tiny\left(C(2-8x+4x^2)+D(6x-12x^2-4x^3)\right)e^{-2x}+4\left(2C(x-x^2)+D(3x^2-2x^3)\right)e^{-2x}+4(C+Dx)x^2e^{-2x}=16xe^{-2x}\]

OpenStudy (anonymous):

lol that ultra small text @_@ \[2Ce^{-2x}+6Dxe^{-2x} = 16xe^{-2x}\]

OpenStudy (unklerhaukus):

how did you simplify that so quick?

OpenStudy (anonymous):

i.. wrote it down on paper, way faster than doing it in head or on comp lol

OpenStudy (unklerhaukus):

that has got to be an easier way, there should be some way to know one of the constants is going to be zero somehow,,

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+4y_p^\prime+4y_p=16xe^{-2x}\] \[y_p=(C+Dx)x^2e^{-2x}\] \[y_p^\prime= (Dx^2+2(C+Dx)x-2(C+Dx)x^2)e^{-2x}\]\[\qquad= (2Cx-2Cx^2+3Dx^2-2Dx^3)e^{-2x}\]\[\qquad\qquad= \left(2C(x-x^2)+D(3x^2-2x^3)\right)e^{-2x}\] \[y_p^{\prime\prime}=\left(2C(1-2x)+D(6x-6x^2)-2\left(2C(x-x^2)+D(3x^2-2x^3)\right)\right)e^{-2x}\]\[=\left(2C-4Cx+6Dx-6Dx^2-4Cx+4Cx^2+6Dx^2-4Dx^3\right)e^{-2x}\]\[=\left(2C-8Cx+4Cx^2+6Dx-12Dx^2-4Dx^3\right)e^{-2x}\]\[=\left(C(2-8x+4x^2)+D(6x-12x^2-4x^3)\right)e^{-2x}\] \[\begin{array}{ccc} \left(C(2-8x+4x^2)+D(6x-12x^2-4x^3)\right)e^{-2x} \\ +4\left(2C(x-x^2)+D(3x^2-2x^3)\right)e^{-2x} \\ +4(C+Dx)x^2e^{-2x} \end{array}=16xe^{-2x} \] \[\left(\begin{array}{ccc} C\left[(2-8x+4x^2)+8(x-x^2)+4x^2\right]\\ +D\left[(6x-12x^2-4x^3)+4\left(3x^2-2x^3\right)+4x^3\right] \end{array}\right) e^{-2x}=16xe^{-2x}\] \[\begin{array}{ccc}C\left[2-8x+4x^2+8x-8x^2+4x^2\right]\\+D\left[6x-12x^2-4x^3+12x^2-8x^3+4x^3\right]\end{array}=16x\] \[2C+6Dx=16x\] \[C=0;\qquad D=\frac 83\] \[y_p=\frac83x^3e^{-2x}\]

OpenStudy (unklerhaukus):

\[y=y_c+y_p\] \[y(x)=(A+Bx)e^{-2x}+\frac83x^3e^{-2x}\]\[y(x)=(A+Bx+\frac83x^3)e^{-2x}\]

OpenStudy (unklerhaukus):

surely there must be a more efficient way at finding the particular solution ,